Im stuck on how to solve the linear system on this problem.
y=6x-11
-2x-3y=-7

Answer by stanbon(75887)

solve the linear system on this problem.
y=6x-11
-2x-3y=-7
-----
Substitute for "y" and solve for "x":
-2x - 3(6x-11) = -7
-2x -18x+33 = -7
-20x = -40
x = 2
----
Solve for "y":
y = 6x - 11
y = 6*2-11
y = 1
=============
Solution: (2,1)
=====================
Cheers,
Stan H.
=====================